# Words Matter: Leveraging Individual Text Embeddings for Code Generation in CLIP Test-Time Adaptation

### v2

@@ Abstract
Vision-language foundation models, such as CLIP, have shown unprecedented zero-shot performance across a wide range of tasks. Nevertheless, these models may be unreliable under distributional shifts, as their performance is significantly degraded. In this work, we explore how to efficiently leverage class text information to mitigate these distribution drifts encountered by large pre-trained vision-language models (VLMs) during test-time inference. In particular, we propose to generate pseudo-labels for the test-time samples by exploiting generic class text embeddings as fixed centroids of a label assignment problem, which is efficiently solved with Optimal Transport. Furthermore, the proposed adaptation method (CLIP-OT) integrates a multiple template knowledge distillation approach, which replicates multi-view contrastive learning strategies in unsupervised representation learning but without incurring additional computational complexity. Extensive experiments on multiple popular test-time adaptation benchmarks presenting diverse complexity empirically show the superiority of CLIP-OT, achieving performance gains of up to $7\%$ over recent state-of-the-art methods, yet being computationally and memory efficient [^1] Code available at https://github.com/ShambhaviCodes/CLIPOT.
@@ 1 Introduction
Large pre-trained vision-language models (VLMs), such as CLIP [33] and ALIGN [19], have emerged as a new paradigm shift in machine learning, revealing promising zero-shot transferability. Nevertheless, if the model is exposed to domain drifts between the training and test samples, its performance can be largely degraded [44, 37].
@@
While a straightforward solution to bridge this gap consists in fine-tuning the trained model using domain-specific labeled data [22, 13], this strategy brings several limitations in real-world scenarios, which may hamper their scalability and deployment. First, adapting to new domains requires collecting labeled samples drawn from each distinct distribution. This might be impractical for specific domains and further hinder the real-time adaptation of the trained model for each input test sample. Furthermore, fine-tuning the model may undermine its desirable zero-shot capabilities [40].
@@
Test-Time Adaptation (TTA) presents a realistic and practical scenario of unsupervised domain adaptation, where a pre-trained model requires real-time adaptation to new data to address unknown distribution drifts without access to supervisory signals [39, 17, 45]. Nevertheless, despite the recent rise of VLMs, and the popularity of TTA in more traditional deep models, such as CNNs and ViTs, the study of TTA in large pre-trained vision-language models remains less explored. Initial approaches, such as TENT [39], which simply minimizes a Shannon entropy objective, have been adopted in CLIP test-time adaptation [36]. In contrast, more recent strategies resort to pseudo-labels on the inference batch to guide the model adaptation [31]. In particular, the core idea of these later methods is to minimize a classification loss, typically the standard cross-entropy, between the generated pseudo-labels and model predictions, which guides the network updates over multiple iterations. Nonetheless, while this strategy is common in semi-supervised learning, where labeled instances are available for all categories, applying it naively in unsupervised scenarios, such as test-time adaptation, can lead to degenerate solutions [17].
In light of the above limitation, we propose a novel approach for CLIP test-time adaptation based on the simple observation that this task can draw inspiration from clustering. Nevertheless, instead of jointly learning the cluster centroids during training, these can be replaced by the fixed class prototypes derived from the text representations. By doing this, the unlabeled samples in a batch will be assigned to the most similar class prototypes, guided by the text embeddings, which serve as a strong class guidance without requiring explicit sample-wise supervision. Furthermore, as our primary goal is to match distributions between different modalities, i.e., class text embeddings and features from the test images, traditional prototype-based clustering methods may lead to suboptimal performance, as they typically assume an unimodal distribution for each cluster. In contrast, we frame the label assignment problem as an Optimal Transport problem, as it can more naturally model multi-modal distributions [23], and is efficiently solved using the Sinkhorn algorithm [9].
Our key contributions can be summarized as follows:
@@
  • We propose to cast the pseudo-labeling strategy in CLIP test-time adaptation as an Optimal Transport assignment, which leverages the class text information available in vision-language models in the form of fixed cluster centroids without requiring further annotations.
@@
  • We introduce a multi-template knowledge distillation approach, which leverages richer information derived from different text prompts to better guide the adaptation without incurring significant computational overhead.
  • Experiments across 240 scenarios demonstrate the superiority of CLIP-OT over recent state-of-the-art methods, all while avoiding additional computational complexity.
@@ 2 Related Work
Test-Time Adaptation.  TTA adapts a pre-trained model to a stream of incoming unlabeled target domain data, processed in batches during testing [39, 47, 8, 30]. Existing approaches for TTA in more traditional deep models can be roughly categorized into: i) normalization-based methods, which leverage the statistics of the test data to adjust the BatchNorm statistics of the model [27, 35]; ii) entropy-based approaches [39, 29, 12], where the model is adapted optimizing the Shannon entropy of the predictions; and iii) pseudo-label strategies [25], which employ the test-time generated labels for supervising the model. With the advent of VLMs, several works have attempted to accommodate some of these techniques to adapt large pre-trained foundation models, particularly CLIP [36, 26, 31, 14], which mainly differs on the different parameter groups updated during adaptation, i.e., prompt tuning [36, 26, 10] or layer-norm [14, 31] strategies. Test-Time Prompt Tuning (TPT) [36] fine-tunes only the input text prompts, which are optimized by minimizing the entropy of the model predictions. To achieve this, TPT generates multiple random augmented views for each test image, incurring additional costs due to performing multiple forward passes through the CLIP vision encoder for each image. Furthermore, TPT is a unimodal approach, which restricts a multi-modal model such as CLIP from fully exploiting its multi-modal capabilities for adaptation. Vision-Text-Space Ensemble (VTE) [10], uses an ensemble of different prompts as input to the CLIP text encoder, where both text and visual encoders remain frozen. Nonetheless, keeping the vision encoder frozen makes it difficult for the model to adapt to images with severe noise effectively. Most recently, CLIPArTT [14] and WATT [31] propose to fine-tune the normalization layers instead, resorting to pseudo-labels computed from image-to-image and text-to-text similarities.
@@
Clustering for unsupervised representation learning.  Jointly learning how to adapt a set of learnable parameters of a deep network and identify the class assignments can be viewed as clustering and unsupervised representation learning. Thus, our work is closely related to recent literature on deep clustering-based approaches [1, 3, 4, 16, 42, 43, 18, 5]. In [3], $k$-means assignments are leveraged as pseudo-labels to learn visual representations, strategy that is later employed in [4] to pre-train standard supervised deep models. Nevertheless, applying naive $k$-means risks collapsing to only a few imbalanced clusters, making it ineffective. The label assignment task is framed as an instance of the Optimal Transport problem in [1], which stresses that the formulation presented in [3] is not principled, since potential degenerate solutions are avoided via specific implementation choices. Caron et al. [5] propose to enforce consistency between cluster assignments produced for different image views, or augmentations, which avoids expensive pairwise comparisons typically performed in contrastive learning. While we consider a similar formulation to some of these works, several vital differences exist. In particular, all these works, e.g., [4, 1, 3] are used in unsupervised representation learning, and thus the prototypes are learned from the data distribution. In contrast, our approach leverages the text representations to enhance the cluster assignments. Note that text embeddings are accessible for free in VLMs, not incurring additional supervision, as the test image label remains unknown. A natural benefit from this strategy is that we do not need to resort to additional augmentation or “multi-views” strategies [5], which introduce a computational burden. Indeed, we can leverage individual text templates as “augmented views”, enhancing the representation learning of the adapted model without incurring additional costs. In fact, these embeddings are computed only once offline. In contrast, image augmentation or “multi-views” strategies must create the additional images for each incoming data batch, which means they perform one forward pass per new augmentation at test time.
@@ 3.1 Vision-language models
[…]
@@
where $\tau$ is a temperature parameter, whose value is learned during pre-training [33], and $\mathbf{z}^{\top}\mathbf{t}$ is the dot product operator.
@@ 3.2 Test-Time Adaptation using CLIP
Vanilla entropy minimization.  Inspired by the semi-supervised learning literature, a straightforward solution to leverage the predicted probability in Eq. 1 in TTA would be to adapt the model parameters based on an entropy minimization objective, similar to TENT [39]:
$\mathcal{L}(p)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}p(y=k|\mathbf{x}_{i})\log p(y=k|\mathbf{x}_{i}),$ (2)
with $N$ denoting the size of the test set, or batch. However, relying on the Shannon entropy [39] in the fully unsupervised case poses a significant risk, as it may potentially result in a degenerate solution, i.e., Eq. (2) might be trivially minimized by assigning all data points to a single, arbitrary label.
4 Proposed solution
The proposed approach aims to overcome the limitations in Eq. 2 objective by finding reliable pseudo-labels in a batch of test-time samples without supervision. To achieve this, we propose CLIP-OT. This novel TTA method aims to better leverage the text embeddings generated by the frozen CLIP text encoder, serving as strong class descriptors to yield label assignments for test samples through Optimal Transport (OT). Note that our motivation drastically differs from the existing self-supervised representation learning literature that also resorts to pseudolabels [3], where cluster centroids are iteratively updated based on the prior label assignments. Below, we elaborate on the details of our method, which is illustrated in Figure 1.
Learning objective.  To overcome the limitation of naively minimizing Eq. 2 in the fully unlabeled scenario, we first encode the model predictions as posterior distributions $q(y|\mathbf{x}_{i})$, which results in:
$\mathcal{L}(p,q)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{K}q(y=k|\mathbf{x}_{i})\log p(y=k|\mathbf{x}_{i}).$ (3)
4.1 Online pseudolabels assignments
@@
$\mathcal{L}(p,q)=-\frac{1}{N}\sum_{i=1}^{N}\left[\frac{1}{\tau}\mathbf{z}_{i}^{\top}\mathbf{T}\mathbf{q}_{i}-\log\sum_{k=1}^{K}\exp(\frac{\mathbf{z}_{i}^{\top}\mathbf{t}_{k}}{\tau})\right],$ (4)
where $\mathbf{T}=[\mathbf{t}_{1},...,\mathbf{t}_{K}]$ represents the matrix containing the class text prototypes, and the logarithmic term in the right side does not depend on $\mathbf{q}$. If we now express Eq. 4 in its matrix form over all the test images, $\mathbf{Z}$, and let $\mathbf{Q}$ be a matrix with columns $\mathbf{q}_{i}$, we have the following objective:
$\max_{\mathbf{Q}\in\mathcal{Q}}\;tr(\mathbf{Q}^{\top}\mathbf{T}^{\top}\mathbf{Z}),$ (5)
where, following [9], we enforce the matrix $\mathbf{Q}$ to be an element of the transportation polytope:
$\mathcal{Q}\mathrel{:}=\{\mathbf{Q}\in^{K\times N}_{+}\mid\mathbf{Q}\mathbf{1}_{N}=\frac{1}{K}\mathbf{1}_{K},\mathbf{Q}^{\top}\mathbf{1}_{K}=\frac{1}{N}\mathbf{1}_{N}\},$ (6)
with $\mathbf{1}_{K}$ and $\mathbf{1}_{N}$ denoting the vectors of ones in dimension $K$ and $N$, respectively. The constraints on $\mathcal{Q}$ enforce that on average each prototype is selected at least $\frac{N}{K}$ times, encouraging $\mathbf{Q}$ to be a matrix with uniform marginals in both rows and columns. Note that such constraints “break” the direct dependency between the original prediction $p$ in Eq. 2 and the posterior $q$ in Eq. 4, since the latter consider the whole batch distribution and its marginal properties for the assignment. This results in more robust supervisory signals, later demonstrated empirically (Figure 2).
Figure 1: CLIP-OT. Our method leverages Optimal Transport (with the Sinkhorn algorithm) to yield soft-codes $Q^{*}_{m}$ (Eq. 9). Then, it minimizes the pseudo cross-entropy between $Q^{*}_{m}$ and the CLIP predictions $P$ as unsupervised loss during test-time adaptation (Eq. 3). Note that, at each test batch, our model runs for $m$ iterations, each leveraging a different text template $m$ to obtain the pseudo-codes $Q^{*}_{m}$.
Codes optimization.  As the objective function in Eq. (5) is linear, and the constraints defining $\mathcal{Q}$ are also linear, this is a linear program. Nevertheless, directly optimizing the above learning objective might be time-consuming, particularly as the number of data points and classes increases. To address this issue and allow a faster optimization, we apply the Sinkhorn algorithm [9], which integrates an entropic constraint, enforcing a simple structure on the optimal regularized transport. Hence, the optimization problem becomes:
@@
where $\mathcal{H}(\mathbf{Q})$ denotes the Shannon entropy function $\mathcal{H}(\mathbf{Q})=-\sum_{ij}\mathbf{Q}_{ij}\log\mathbf{Q}_{ij}$ and $\varepsilon$ controls its weight. Furthermore, to align with TTA assumptions, we work on batches by restricting the transportation polytope to the current batch [5], in contrast to other works that employ the full dataset [1]. Thus, the dimensionality of the $\mathbf{Q}$ matrix becomes $K\times B_{s}$, where $B_{s}$ denotes the batch size. Now, the soft codes $\mathbf{Q}^{*}$ are the solution of the problem presented in (7) over the set $\mathcal{Q}$, which can be efficiently optimized with a few iterations as:
@@
where $\mathbf{u}$ and $\mathbf{v}$ are renormalization vectors in K and B respectively, with $t$ denoting the iteration.
Optimizing codes for individual textual views.  Very recent literature [31] has pointed out the limitations of using the average text prototypes, and has instead suggested to resort to multiple category embeddings, each obtained through the different text templates in $\mathcal{T}$. Thus, we modify the solution presented in Eq. 8 to be specialized for specific class template [^2] The text templates are the same as in WATT [31] – see Appendix A., where the codes for each template $m$ are:

### v3

# Semantic Anchor Transport: Robust Test-Time Adaptation for Vision-Language Models

@@ Abstract
Large pre-trained vision-language models (VLMs), such as CLIP, have shown unprecedented zero-shot performance across a wide range of tasks. Nevertheless, these models may be unreliable under distributional shifts, as their performance is significantly degraded. In this work, we investigate how to efficiently utilize class text information to mitigate distribution drifts encountered by VLMs during inference. In particular, we propose generating pseudo-labels for the noisy test-time samples by aligning visual embeddings with reliable, text-based semantic anchors. Specifically, to maintain the regular structure of the dataset properly, we formulate the problem as a batch-wise label assignment, which is efficiently solved using Optimal Transport. Our method, Semantic Anchor Transport (SAT), utilizes such pseudo-labels as supervisory signals for test-time adaptation, yielding a principled cross-modal alignment solution. Moreover, SAT further leverages heterogeneous textual clues, with a multi-template distillation approach that replicates multi-view contrastive learning strategies in unsupervised representation learning without incurring additional computational complexity. Extensive experiments on multiple popular test-time adaptation benchmarks presenting diverse complexity empirically show the superiority of SAT, achieving consistent performance gains over recent state-of-the-art methods, yet being computationally efficient.
@@ 1 Introduction
Large pre-trained vision-language models (VLMs), such as CLIP [radford2021learning] and ALIGN [jia2021scaling], have emerged as a new paradigm shift in machine learning, revealing promising zero-shot transferability. Nevertheless, if the model is exposed to domain drifts at test time, its performance can be largely degraded [yu2023task, silva2024closer]. While a straightforward solution to bridge this gap involves fine-tuning the trained model using domain-specific labeled data [lai2023padclip, goyal2023finetune], this strategy presents several limitations in real-world scenarios, which may hinder its scalability and deployment. First, adapting to new domains requires collecting labeled samples drawn from each distinct distribution. This might be impractical for specific domains and further hinder the real-time adaptation of the trained model for each input test sample. Furthermore, fine-tuning the model may undermine its desirable zero-shot capabilities [WiSE].
@@
Test-Time Adaptation (TTA) presents a realistic and practical scenario for unsupervised domain adaptation, where a pre-trained model requires real-time adaptation to new data to address unknown distribution drifts without access to supervisory signals [wangtent, iwasawa2021test, yuan2023robust]. Nevertheless, despite the recent rise of VLMs, and the popularity of TTA in more traditional deep models, such as CNNs and ViTs, the study of TTA in large pre-trained vision-language models remains less explored. Standard approaches, e.g., TENT [wangtent], which minimizes a Shannon entropy objective, have been adopted in CLIP test-time adaptation [shu2022test]. On the other hand, more recent strategies utilize pseudo-labels within the inference batch to guide model adaptation [osowiechi2024watt, hakim2024clipartt, maharana2024texttt]. In particular, the core idea of these later methods is to minimize a classification loss, typically the standard cross-entropy, between the generated pseudo-labels and model predictions, which guides the network updates over multiple iterations. Nonetheless, while this strategy is common in semi-supervised learning, where additional labeled instances are available for all categories, applying it naively in unsupervised scenarios, such as test-time adaptation, can lead to degenerate solutions over multiple updates [iwasawa2021test], a problem we refer to as error accumulation.
Figure 1: Error Accumulation. We track examples from all corruptions from CIFAR10C dataset with initial zero-shot CLIP predictions (at $t=0$) that are misclassified. Baselines [wangtent, maharana2024texttt, osowiechi2024watt] catastrophically reinforce this error: similarity to the ‘Wrong Class’ (dashed lines) increases while similarity to the ‘True Class’ (solid lines) decreases. In contrast, our method, SAT (orange), is the only one that provides a corrective signal, actively reducing similarity to the wrong class and breaking the cycle of error accumulation.
This problem is particularly prevalent in the literature on test-time adaptation in VLMs. To illustrate this phenomenon, we conducted a simple experiment, the findings of which are depicted in Figure 1. Considering all corruptions in CIFAR10C, we identify samples misclassified by zero-shot CLIP and track their behavior across consecutive TTA steps. Specifically, we compute the average cosine similarity between the adapted visual embeddings and the corresponding class text embeddings of both the wrong (i.e., predicted) and true (i.e., expected) category. Ideally, a robust TTA method should reduce the similarity to the wrong class prototype, gradually correcting the initial misalignment. Nevertheless, as shown in Figure 1, existing methods tend to reinforce these mistakes, thereby increasing the similarity between the visual and text prototypes of the incorrect class, even when the prediction is wrong.
In light of the above limitation, we aim to provide a more robust TTA framework that can correct incorrect initial supervisory signals. To do so, we reframe TTA as a cross-modal alignment problem, where the goal is to align the embeddings of test images to their corresponding fixed semantic anchors provided by the language modality. This alignment can be viewed through the lens of clustering, where the unlabeled samples are grouped around class text prototypes. Hence, we can take into account the structure of the joint test batch to better correct sample-specific inconsistencies. Nevertheless, instead of jointly learning the cluster centroids during training, we use the fixed class prototypes derived from the text representations, which provide stable semantic guidance without requiring explicit supervision. A key challenge is the distribution mismatch between text prototypes and test image embeddings, since they originate from different modalities. Traditional clustering methods, which typically assume unimodal distributions, often struggle to bridge this gap, resulting in suboptimal performance. In contrast, we formulate label assignment as an Optimal Transport (OT) problem, yielding a global, cost-aware correspondence between visual embeddings and text prototypes that naturally model multi-modal distributions [multimodal_dist_ot]. This perspective is more robust to outliers or noisy predictions, directly mitigating the error accumulation that undermines existing self-training in TTA.
The main contributions of this paper can be summarized as:
@@
  • We propose Semantic Anchor Transport (SAT). This novel framework casts the pseudo-labeling strategy in CLIP test-time adaptation as a batch-wise Optimal Transport assignment, which leverages the class text information available in vision-language models in the form of fixed robust cluster centroids without requiring further annotations.
@@
  • We introduce a multi-template knowledge distillation approach that leverages richer information derived from different text prompts to better guide adaptation without incurring significant computational overhead.
  • Comprehensive experiments across multiple visual corruptions and domain shifts benchmarks demonstrate the superiority of SAT over recent state-of-the-art methods.
@@ 2 Related Work
Test-Time Adaptation (TTA) aims at adapting a pre-trained model to a stream of incoming unlabeled target domain data, processed in batches during testing [wangtent, zhang2022memo, choi2022improving, niutowards]. Existing approaches in traditional unimodal models can be roughly categorized into: i) normalization-based methods, which leverage the statistics of the test data to adjust the BatchNorm statistics of the model [mirza2022norm, schneider2020improving]; ii) entropy-based approaches [wangtent, niu2022efficient, goyal2022test], where the model is adapted optimizing the Shannon entropy of the predictions; and iii) pseudo-label strategies [liang2020we], which employ the test-time generated labels for supervising the model. With the advent of VLMs, several works have attempted to accommodate some of these techniques to adapt pre-trained foundation models, notably CLIP [shu2022test, ma2024swapprompt, osowiechi2024watt, hakim2024clipartt], which mainly differ on the different parameter groups updated during adaptation, i.e., prompt tuning [shu2022test, ma2024swapprompt, dobler2024lost] or layer-norm [hakim2024clipartt, osowiechi2024watt] strategies. For example, Test-Time Prompt Tuning (TPT) [shu2022test] optimizes input text prompts by minimizing the model’s prediction entropy. In contrast, Vision-Text-Space Ensemble (VTE) [dobler2024lost] uses an ensemble of prompts as input to the text encoder. Nevertheless, keeping both the text and vision encoder frozen makes it difficult for the model to adapt to images with severe noise effectively. Recent VLM adaptation methods, such as WATT [osowiechi2024watt] and BATCLIP [maharana2024texttt], which employ parameter-efficient updates of the encoders, have developed sophisticated bimodal heuristics to improve pseudo-label quality. WATT [osowiechi2024watt] generates them by combining both image-to-image and text-to-text similarity matrices. Taking this a step further, BATCLIP [maharana2024texttt] first generates pseudo-labels via a standard argmax prediction, which are refined via additional losses. Nevertheless, both approaches rely on locally derived pseudo-labels, failing to explicitly optimize the sample-to-class assignment group cost, making them more sensitive to outliers or ambiguous samples.
@@
Clustering for unsupervised representation learning. Jointly adapting the parameters of a deep network while inferring the class assignments can be viewed as clustering and unsupervised representation learning. Thus, our work is closely related to recent literature on deep clustering-based approaches [asanoself, caron2018deep, caron2019unsupervised, huang2019unsupervised, yan2020clusterfit, yang2016joint, jabi2019deep, caron2020unsupervised]. In [caron2018deep], $k$-means assignments are leveraged as pseudo-labels to learn visual representations, a strategy later employed in [caron2019unsupervised] to pre-train standard supervised deep models. Nevertheless, applying naive $k$-means risks collapsing to only a few imbalanced clusters, making it ineffective. A more principled approach consists in framing the label assignment task as an instance of the Optimal Transport problem [asanoself], whose global, batch-aware optimization and balancing constraints naturally prevent degenerate solutions. Furthermore, Caron et al. [caron2020unsupervised] enforce consistency between cluster assignments across different image views or augmentations, avoiding expensive pairwise comparisons typically performed in contrastive learning. While our formulation shares similarities, key differences exist. In particular, all these methods, i.e., [caron2019unsupervised, asanoself, caron2018deep] operate in unsupervised representation learning, where prototypes are learned from the data distribution. In contrast, we leverage the text representations as fixed semantic anchors. Note that text embeddings are accessible for free in VLMs, not incurring additional supervision, as the test image label remains unknown. A natural benefit from this strategy is that we do not need to resort to additional augmentation or “multi-views” strategies [caron2020unsupervised], which introduce a computational burden. Instead, we treat individual text templates as “augmented views”, enhancing the representation learning of the adapted model without incurring extra cost. In fact, these embeddings are computed only once offline, whereas image augmentation or “multi-views” strategies must create the additional images for each incoming batch, performing one forward pass per new augmentation at test time.
@@ 3.1 Vision-language models
CLIP [radford2021learning] is a foundation vision-language model, trained via contrastive learning to produce visual representations from images $\mathbf{x}$ paired with their associated text descriptions $T$. To do so, CLIP consists of an image encoder $\boldsymbol{\theta}$ and a text encoder $\boldsymbol{\phi}$. This generates the corresponding vision $\mathbf{z}\in\real^{d}$ and class text $\mathbf{t}_{k}\in\real^{d}$ embeddings (column vectors), which are typically projected into an $\ell_{2}$-normalized shared embedding space. At inference, this learning paradigm enables zero-shot prediction. More concretely, for a given set of $K$ classes, and an ensemble of $M$ different templates, we can generate the set of available prompts as $\mathcal{T}=\{\{T_{mk}\}_{m=1}^{M}\}_{k=1}^{K}$, whose embedding for template $m$ and class $k$ is obtained as $\mathbf{t}_{mk}=\boldsymbol{\phi}(\textsc{{\color[rgb]{.5,.5,.5}\definecolor[named]{pgfstrokecolor}{rgb}{.5,.5,.5}\pgfsys@color@gray@stroke{.5}\pgfsys@color@gray@fill{.5}``A photo of a [class${}_{k}$]''}})$. Then, a popular strategy [radford2021learning, clipad, WiSE] consists in obtaining a class zero-shot prototype, which is computed as $\mathbf{t}_{k}=\frac{1}{M}\sum^{M}_{m=1}\boldsymbol{\phi}(T_{mk})$. Then, for a given test image $\mathbf{x}_{i}$, its zero-shot prediction, $\mathbf{p}_{i}=(p_{ik})_{1\leq k\leq K}$, can be obtained as:
@@
where $\tau$ is a temperature parameter learned during pre-training [radford2021learning], and $\mathbf{z}^{\top}\mathbf{t}$ indicates dot product.
@@ 3.2 Test-Time Adaptation using CLIP
Vanilla entropy minimization. Inspired by the semi-supervised learning literature, a straightforward solution to leverage the predicted probability in Eq. 1 in TTA would be to adapt the model parameters based on an entropy minimization objective, similar to TENT [wangtent]:
$\mathcal{L}(p)=-\frac{1}{B_{T}}\sum_{i=1}^{B_{T}}\sum_{k=1}^{K}p(y=k|\mathbf{x}_{i})\log p(y=k|\mathbf{x}_{i}),$ (2)
with $B_{T}$ denoting the size of the test batch. However, relying on the Shannon entropy [wangtent] in the fully unsupervised case poses a significant risk, as it may potentially result in a degenerate solution, i.e., Eq. (2) might be trivially minimized by assigning all data points to a single, arbitrary label.
4 SAT: TTA as Cross-Modal Alignment
The proposed approach aims to overcome the limitations in Eq. 2 objective by finding reliable pseudo-labels in a batch of test-time samples without supervision. To achieve this, we propose SAT. We formulate TTA as a cross-modal alignment problem, where the core task is to generate a robust supervisory signal for each unlabeled test batch. This novel TTA method leverages the text embeddings generated by the frozen CLIP text encoder more effectively, serving as strong class descriptors to yield label assignments for test samples, which we propose propagating through Optimal Transport. Note that our motivation drastically differs from the existing self-supervised representation learning literature that also resorts to pseudolabels [caron2018deep], where cluster centroids are iteratively updated based on the prior label assignments. Furthermore, by solving a globally optimal matching problem, SAT is more robust to noisy predictions than existing TTA methods [osowiechi2024watt, maharana2024texttt]. Below, we elaborate on the details of our method, which is illustrated in Figure 2.
Figure 2: SAT leverages Optimal Transport (with the Sinkhorn algorithm) to yield soft-codes $Q^{*}_{m}$ (Eq. 9). Then, it minimizes the pseudo cross-entropy between $Q^{*}_{m}$ and the CLIP predictions $P$ as unsupervised loss during test-time adaptation (Eq. 3). Note that, at each test batch, our model runs for $m$ iterations, each using a different text template $m$ to obtain the pseudo-codes $Q^{*}_{m}$.
To overcome the limitation of naively minimizing Eq. 2 in the fully unlabeled scenario, TTA literature typically encodes the model predictions as posterior distributions $q(y|\mathbf{x}_{i})$, which results in:
$\mathcal{L}(p,q)=-\frac{1}{B_{T}}\sum_{i=1}^{B_{T}}\sum_{k=1}^{K}q(y=k|\mathbf{x}_{i})\log p(y=k|\mathbf{x}_{i}).$ (3)
Thus, the adaptation is driven by the pseudo cross-entropy in Eq. (3), where a significant challenge lies in constructing a high-quality target distribution, or pseudo-label $q_{ik}=q(y=k|\mathbf{x}_{i})$, without access to ground truth labels. Indeed, as illustrated in Figure 1, prior TTA methods tend to amplify prediction errors, leading to a state of error accumulation. Our central thesis is that we can break this loop by exploiting the natural alignment between visual features and their text-based semantic anchors in the batch.
4.1 Batch-Aware Cross-modal Alignment
@@
$\mathcal{L}(p,q)=-\frac{1}{B_{T}}\sum_{i=1}^{B_{T}}\left[\frac{1}{\tau}\mathbf{z}_{i}^{\top}\mathbf{T}\mathbf{q}_{i}-\log\sum_{k=1}^{K}\exp(\frac{\mathbf{z}_{i}^{\top}\mathbf{t}_{k}}{\tau})\right],$ (4)
where $\mathbf{T}=[\mathbf{t}_{1},...,\mathbf{t}_{K}]$ represents the matrix containing the class text prototypes, and the logarithmic term in the right side does not depend on $\mathbf{q}$. If we now express Eq. 4 in its matrix form over all the test images in the batch (with $\mathbf{Z}\in\mathbb{R}^{d\times B_{T}}$ the corresponding embeddings), and let $\mathbf{Q}$ be a matrix with columns $\mathbf{q}_{i}$, we have the following objective:
$\max_{\mathbf{Q}\in\mathcal{Q}}\;tr(\mathbf{Q}^{\top}\mathbf{T}^{\top}\mathbf{Z}).$ (5)
which seeks the optimal alignment between the batch’s visual features $\mathbf{Z}\in\mathbb{R}^{d\times B_{T}}$ and a set of text-based “semantic anchors” $\mathbf{T}\in\mathbb{R}^{d\times K}$. Following [cuturi2013sinkhorn], we enforce the matrix $\mathbf{Q}$ to be an element of the transportation polytope:
$\mathcal{Q}\mathrel{:}=\{\mathbf{Q}\in\real^{K\times B_{T}}_{+}\mid\mathbf{Q}\mathbf{1}_{B_{T}}=\frac{1}{K}\mathbf{1}_{K},\mathbf{Q}^{\top}\mathbf{1}_{K}=\frac{1}{B_{T}}\mathbf{1}_{B_{T}}\},$ (6)
with $\mathbf{1}_{K}$ and $\mathbf{1}_{N}$ denoting the vectors of ones in dimension $K$ and $N$, respectively. The constraints on $\mathcal{Q}$ enforce that on average each prototype is selected at least $\frac{N}{K}$ times, encouraging $\mathbf{Q}$ to be a matrix with uniform marginals in both rows and columns. Note that such constraints ‘break’ the direct dependency between the original prediction $p$ in Eq. 2 and the posterior $q$ in Eq. 4, since the latter considers the whole batch distribution and its marginal properties for the assignment. This results in more robust supervisory signals, later demonstrated empirically (Figure 3).
Finding optimal Q. As the objective function in Eq. (5) is linear, and the constraints defining $\mathcal{Q}$ are also linear, this is a linear program. Nevertheless, directly optimizing the above learning objective might be time-consuming, particularly as the number of data points and classes increases. To address this issue and facilitate faster optimization, we apply the Sinkhorn algorithm [cuturi2013sinkhorn], which incorporates an entropic constraint that enforces a simple structure on the optimal regularized transport. Hence, the optimization problem becomes:
@@
where $\mathcal{H}(\mathbf{Q})$ denotes the Shannon entropy function $\mathcal{H}(\mathbf{Q})=-\sum_{ij}\mathbf{Q}_{ij}\log\mathbf{Q}_{ij}$ and $\varepsilon$ controls its weight. Furthermore, to align with TTA assumptions, we work on batches by restricting the transportation polytope to the current batch [caron2020unsupervised], in contrast to other works that employ the full dataset [asanoself]. Thus, the dimensionality of the $\mathbf{Q}$ matrix becomes $K\times B_{T}$, where $B_{T}$ denotes the batch size. Now, the soft assignment matrix $\mathbf{Q}^{*}$ is the solution of the problem in (7) over the set $\mathcal{Q}$, which can be efficiently optimized with a few iterations:
@@
with $\mathbf{u}$ and $\mathbf{v}$ representing renormalization vectors in K and $\real{}^{B_{T}}$ respectively, and $t$ the iteration.
Benefits of the proposed strategy. The proposed Optimal Transport (OT) based formulation provides: i) Global consistency, since OT is batch-aware. The assignment for any single image is dependent on the features of all other images in the batch. The solver finds the best overall configuration, making it inherently robust to individual outliers. ii) Intrinsic regularization given by the polytope constraints (Eq. 6), which enforce that, on average, all classes are represented in the pseudo-labels. This structurally prevents the model from collapsing to a single class, yielding robustness.
4.2 Semantic alignment with multiple templates
Recent literature [osowiechi2024watt] has pointed out the limitations of using the average text prototypes, and has instead suggested to resort to multiple category embeddings, each obtained through the different text templates in $\mathcal{T}$. Thus, to create a richer and more robust supervisory signal, we leverage the ensemble of $M$ diverse text templates typically used in TTA, each providing a unique “semantic view”, $\mathbf{T}_{m}=[\mathbf{t}_{1m},...,\mathbf{t}_{Km}]$, where $\mathbf{T}_{m}$ denotes the matrix containing the class text prototypes obtained from the $m$-th template. Compared to existing TTA approaches, we introduce a more sophisticated knowledge distillation strategy that disentangles the goals of assignment and generalization.
1
For Assignment, we need specificity. We compute a separate, “clean” transport plan $\mathbf{Q}^{*}_{m}$ for each specific semantic view $\mathbf{T}_{m}$. This provides an unambiguous matching based on a single, clear context (e.g., aligning to “a sketch of a class” vs. “a photo of a class”). Concretely, we modify the solution presented in Eq. 8 to be specialized for specific class template [^1] The text templates are the same as in WATT [osowiechi2024watt] – see Appendix A., where the codes for each template $m$ are: